If 4 - 3x = 24 + 2x, evaluate 3x2 - 2x + 4
the 3x2 is 3 x to the second power.
Solve 3 + 2[4x - 3(5 - x 0] = 3(x - 20)

Three-fifths of a number is negative thirty. Find the number.
A rectangular vegetable grden has a perimeter of 64 ft. The length of the garden is 20 ft. Find the width of the garden.

Answer by stanbon(75887)

If 4 - 3x = 24 + 2x, evaluate 3x^2 - 2x + 4
-----------------------------
If 4-3x=24+2x
then 5x = -20
and x=-4
-------------
Evaluate 3x^2 - 2x + 4 when x=-4
3(-4)^2-2(-4)+4
=48+8+4
=60
----------
I am going to assume you mean x 0 is x^0'
If that is true your problem becomes:
Solve 3 + 2[4x - 3(5 - x^0] = 3(x - 20)
3+8x-15+3 = 3x-60
5x=-51
x=-51/5
-------------------
Three-fifths of a number is negative thirty. Find the number.
Let the number be "x".
EQUATION:
(3/5)x=-30
Multiply both sides by (5/3) to get:
x=-50
----------------


A rectangular vegetable garden has a perimeter of 64 ft. The length of the garden is 20 ft. Find the width of the garden.
Perimeter = 2length + 2 width
64=2(20) + 2width
24=2width
width=12 ft.
-----
Cheers,
Stan H.